Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
cifaldi - 0:08 Monday 20 November 2023 (62498) Print this report
Restoring of the TCS configuration

After the afternoon shift, I restored the TCS configuration settled on November, 17th 2023. 

At 21:57 UTC PR CHRoCC VDAC 0.42 --> 0.08 (21600 sec ramp time) by the operator (see entry 62497). 

At 22.28 UTC SR RH switched off via telnet. ( I try to use the VPM command, but it seems not to respond). 

The settled CO2 actuator's powers are listed in the table below: 

 

Sun Nov 19th 

22.50 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.036

0

0

on the pickoff

0.225

0

0

N

on the ITF

0.024

0.08 

0.55 

on the pickoff

0.147 

0.013 

0.089 

The checks have been performed between 22.32 UTC and 22.55 UTC.

In the end, all the CO2 actuators needed for CITF and lock acquisition flip mirrors have been opened.

Comments to this report:
taranto - 10:21 Monday 20 November 2023 (62500) Print this report

After being called by the control room, at 8.30 UTC, the CO2 powers were checked. The NI DAS IN and OUT showed no issues, the WI CH was changed by 2%, while the NI CH by 14% in order to restore the CO2 actuators powers needed for CITF and lock acquisition. Moreover, at 09.03 UTC PR CHRoCC VDAC 0.08 --> 0.12 (3600 sec ramp time) by the operator.

Search Help
×

Warning

×